Grade 9Mechanics

A car runs with negative acceleration of -0.5 m/s2 . The initial velocity of car is 54 km/hr in how much time and how far from initial point will car stop

sughosh deshpande
10 Years agoGrade 9

1 Answer

Vikas TU
10 Years ago
change 54 in m/s first.
u = 54000/3600 = 540/36  = 60/4 = 30/2 = 15 m/s
v = 15 -0.5t
check when v = 0
15 = 0.5t
t = 150/5 = 30s
at 30 s it will stop.
 
from 2nd law of eqn.
s = ut^2 – 0.5at^2
s = 15*900 – 0.5*0.5*300
s = 100(135 – 0.75)
s = 100*134.25
s = 13425 m
or
s = 13.425 km